OpenStudy (anonymous):

Calculate the wavelength of radiation with a frequency of 1.50 x 10^13Hz ?

OpenStudy (anonymous):

wavelength=1/frequency

OpenStudy (jamesj):

@dockworker. Not quite. http://openstudy.com/study?login#/updates/4f440548e4b065f388dce878

OpenStudy (anonymous):

\[v = \lambda f\]\[\lambda = \frac{v}{f}\]You are given the frequency, and because it's radiation, you know the velocity is the speed of light:\[\lambda = \frac{2.998 \times 10^{8}\frac{m}{s}}{1.5 \times 10^{13}s^{-1}} = 1.998 \times 10^{-5}m\]